# Variational quantum state preparation for quantum-enhanced metrology in noisy systems

### v1

@@ IV Conclusion
There are a number of interesting future directions that can build on the analysis presented in this work. First, we can expand the circuit ansatz in Fig. 1. For example, we can increase the depth of our VQC by layering alternating rotations and entangling stages. This will, in principle, increase the complexity and range of potential states that we can prepare, potentially modifying our current classification into perturbed, squeezed-like, and uncorrelated states. The VQC ansatz can also be expanded to fully exploit the programmability of quantum sensing hardware by including different site-resolved qubit rotations and two-qubit gates, as opposed to global rotations and entangling operations.
@@
Second, in addition to decoherence introduced during the entangling stage of the VQC, we can incorporate decoherence introduced by errors in VQC parameters such as imperfect initial and final rotations (i.e., shot-to-shot fluctuations in rotation axis or angle). In contrast to the current VQC that we study, this could benefit theoretical understanding by introducing some degree of control over the level of mixedness of the prepared states that is independent of the VQC parameters (whereas in the current protocol the total amount of decoherence is correlated with the duration of the entangling stage).
@@
Finally, we are investigating the development of more sophisticated, specialized numerical optimization routines for optimizing the QFI for a given VQC or state preparation protocol. Whereas general-purpose optimization methods, such as the Nelder-Mead algorithm used to generate the results in this manuscript, are designed to receive only scalar outputs, the QFI form in Eq. (6) has considerable structure that can be exploited. Instead of receiving only the scalar value $F_{Q}[\hat{\rho}(\mathbf{x});\hat{G}]$, we are developing optimization methods that can receive and work with the density matrix $\hat{\rho}(\mathbf{x})$ and approximate (finite-differences) or exact (algorithmic-differentiated) derivatives for entries of $\hat{\rho}(\mathbf{x})$ with respect to the parameters $\mathbf{x}$. It is not straightforward to pass this information through an eigendecompostion, which is needed in Eq. (6). However, by building on past work by some of us Larson et al. (2021); Larson and Menickelly (2024) that optimizes composite objectives by using objective knowledge form (such as the form of Eq. (6)) in a QFI-only numerical optimization routine, we hope to find higher-quality solutions in fewer queries/simulations of the system. This is a particularly important task, as simulation of open quantum systems involving many qubits and subsequent computation of the QFI for a given mixed quantum state is computationally expensive.
V Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research, Exploratory Research for Extreme-Scale Science program.
@@ Appendix A Husimi probability distributions
Here we define the Husimi probability distributions used to visualize spin states in Fig. 3. The basic idea is to define a Husimi probability distribution within each manifold of fixed spin length $S$, and take an appropriately weighted average of these distributions. Note that for ease of language, in this appendix we treat the symbol $S$ as a variable to denote spin length, whereas the main text defines $S=N/2$ for $N$ spins.
@@ A.1 A simplified case
and the normalization factor $w_{S}$ is determined by requiring that $\rho(\bm{v})$ is a normalized probability distribution, with

### v2

@@ V Conclusion
A number of interesting future directions can build on the analysis presented in this work. First, we can expand the circuit ansatz in Fig. 1 by, for example, increasing the depth of our VQC by layering alternating rotations and entangling stages. This will, in principle, increase the complexity and range of potential states that we can prepare, potentially modifying our current classification into cat-like, squeezed-like, and uncorrelated states. The VQC ansatz can also be expanded to fully exploit the programmability of quantum sensing hardware by including different site-resolved qubit rotations and two-qubit gates, as opposed to global rotations and entangling operations.
@@
Second, in addition to decoherence introduced during the entangling stage of the VQC, we can incorporate decoherence introduced by errors in VQC parameters such as imperfect initial and final rotations (i.e., shot-to-shot fluctuations in rotation axis or angle). In contrast to the current VQC that we study, this could benefit theoretical understanding by introducing some degree of control over the level of mixedness of the prepared states that is independent of the VQC parameters (whereas in the current protocol, the total amount of decoherence is correlated with the duration of the entangling stage).
@@
We are investigating the development of more sophisticated, specialized numerical optimization routines for optimizing the QFI for a given VQC or state preparation protocol. Whereas general-purpose optimization methods, such as the Nelder–Mead algorithm used to generate the results in this manuscript, are designed to receive only scalar outputs, the QFI form in Eq. (6) has considerable structure that can be exploited. Instead of receiving only the scalar value $F_{Q}[\hat{\rho}(\mathbf{x});\hat{G}]$, we are developing optimization methods that can receive and work with the density matrix $\hat{\rho}(\mathbf{x})$ and approximate (finite-differences) or exact (algorithmic-differentiated) derivatives for entries of $\hat{\rho}(\mathbf{x})$ with respect to the parameters $\mathbf{x}$. It is not straightforward to pass this information through an eigendecomposition, which is needed in Eq. (6). However, by building on past work by some of us Larson et al. (2021); Larson and Menickelly (2024) that optimizes composite objectives by using objective knowledge form (such as the form of Eq. (6)) in a QFI-only numerical optimization routine, we hope to find higher-quality solutions in fewer queries/simulations of the system. This is a particularly important task because simulation of open quantum systems involving many qubits and subsequent computation of the QFI for a given mixed quantum state are computationally expensive.
Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research, Exploratory Research for Extreme-Scale Science program. The majority of computing for this project was performed at the OU Supercomputing Center for Education & Research (OSCER) at the University of Oklahoma (OU).
@@ Appendix A Husimi probability distributions
Here we define the Husimi probability distributions used to visualize spin states in Fig. 3. The basic idea is to define a Husimi probability distribution within each manifold of fixed spin length $S$ and take an appropriately weighted average of these distributions. Note that, for ease of language, in this appendix we treat the symbol $S$ as a variable to denote spin length, whereas the main text defines $S=N/2$ for $N$ spins.
@@ A.1 A simplified case
and the normalization factor $w_{S}$ is determined by requiring that $\rho(\bm{v})$ be a normalized probability distribution, with